OpenStudy (anonymous):

If you take Coleman's test grade and divide by 2 and add 27, you get 67. Write and solve an equation to find Coleman's test grade.

OpenStudy (anonymous):

x/2 +27=67 x/2=67-27 x=2(67-27)